OpenStudy (anonymous):

Can someone show me how to find the Zeros from this equation x^3-2x^2-x+2

OpenStudy (anonymous):

\[x ^{3}-2x ^{2}-x+2=0\]

OpenStudy (anonymous):

Do you remember factoring in pairs?

OpenStudy (anonymous):

such as \[x ^{3}-x and -2x ^{2}+2\]

OpenStudy (anonymous):

\[x^3-2x^2-x+2=0\\ x^2(x-2)-(x-2)=0\\ (x^2-1)(x-2)=0\\ (x+1)(x-1)(x-2)=0\\ x+1=0\lor x-1=0\lor x-2=0\\ x=-1\lor x=1\lor x=2\]

OpenStudy (anonymous):

ok I got that thank you so much for your helpe!

OpenStudy (anonymous):

Factor: \[x ^{3}-2x ^{2}\] Factor: \[-x+2\]

OpenStudy (anonymous):

By factoring for the last one, I mean: \[-1(x+2)\]